# SDEs with supercritical distributional drifts

### v1

@@ 3.2. Supercritical case: $\frac{d}{p}+\frac{2}{q}<2+\alpha$
By [49, Theorem 2.2], there is a constant $C=C(\Theta,\alpha,p,q)>0$ such that for all $n\in{\mathbb{N}}$,
$\|u_{n}\|_{{\mathbb{L}}^{\infty}_{T}}+\|u_{n}\|_{{\mathscr{V}}_{T}}\lesssim_{C}\|f_{n}\|_{{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}}\lesssim_{C}\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}}.$ (3.14)
By (2.9) and (2.10), it is easy to see that
$\|\partial_{t}u_{n}\|_{{\mathbb{L}}^{2}_{T}{\mathbf{H}}^{-3}_{2}}$ $\leqslant\|\Delta u_{n}+\mathord{{\rm div}}(b_{n}u_{n})-(\mathord{{\rm div}}b_{n})u_{n}+f_{n}\|_{{\mathbb{L}}^{2}_{T}{\mathbf{H}}^{-3}_{2}}$
$\lesssim\|u_{n}\|_{{\mathbb{L}}^{2}_{T}{\mathbf{H}}^{-1}_{2}}+\|b_{n}u_{n}\|_{{\mathbb{L}}^{2}_{T}{\mathbf{H}}^{-2}_{2}}+\|(\mathord{{\rm div}}b_{n})u_{n}\|_{{\mathbb{L}}^{2}_{T}{\mathbf{H}}^{-3}_{2}}+\|f_{n}\|_{{\mathbb{L}}^{2}_{T}{\mathbf{H}}^{-3}_{2}}\leqslant C,$
where $C$ is independent of $n$. This implies that
$\|u_{n}(t)-u_{n}(s)\|_{{\mathbf{H}}^{-3}_{2}}\leqslant\|\partial_{t}u_{n}\|_{{\mathbb{L}}^{2}_{T}{\mathbf{H}}^{-3}_{2}}|\sqrt{|t-s|}\leqslant C\sqrt{|t-s|}.$
By Aubin-Lions’ lemma, there is a function $u\in{\mathbb{L}}^{\infty}_{T}\cap{\mathscr{V}}_{T}$ and subsequence $\{n_{k}\}$ such that for each $t$ and $\varphi\in C^{\infty}_{c}({\mathbb{R}}^{d})$,
@@
$\|\langle(b_{n_{k}}-b)u_{n_{k}},\nabla\varphi\rangle\|_{L^{1}_{T}}\lesssim_{C}\|(b_{n_{k}}-b)\chi\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{H}}^{\alpha_{b}}_{p_{b}}}\|u_{n_{k}}\|_{{\mathscr{V}}_{T}}\|\nabla\varphi\|_{{\mathscr{V}}_{T}}\stackrel{{\scriptstyle k\to\infty}}{{\to}}0,$
@@
Hence, by $\mathord{{\rm div}}b\in L^{\infty}$ and commutating the order of limits (guaranteeing by the above two limits),
$\lim_{n\to\infty}\langle u,b\cdot\nabla u_{n}\rangle=\lim_{n\to\infty}\lim_{m\to\infty}\langle u_{m},b\cdot\nabla u_{n}\rangle$ $=\lim_{n\to\infty}\lim_{m\to\infty}\langle\mathord{{\rm div}}b\,u_{m},u_{n}\rangle-\lim_{n\to\infty}\lim_{m\to\infty}\langle b\cdot\nabla u_{m},u_{n}\rangle$
$=\langle\mathord{{\rm div}}b\,u,u\rangle-\lim_{m\to\infty}\langle b\cdot\nabla u_{m},u\rangle,$
@@
$\lim_{n\to\infty}\langle u,b\cdot\nabla u_{n}\rangle=\langle\mathord{{\rm div}}b\,u,u\rangle/2.$
@@
[…]
@@
  1. $\bullet$
    For any bounded measurable function $g:[0,T]\times{\mathbb{R}}^{d}\to{\mathbb{R}}$ with that the function $x\to g(t,x)$ is uniformly Lipschitz with respect to $t\in[0,T]$, the product $gf$ is well-defined for all $f\in{\mathbb{B}}$, and there is a constant $C_{g}>0$ such that for all $f\in{\mathbb{B}}$,
@@ 1. $\bullet$
$\lim_{n\to\infty}\|f_{n}-f\|^{q}_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q}}\leqslant\int^{T}_{0}\sum_{j\geqslant-1}2^{\alpha jq}\lim_{n\to\infty}\|\phi_{n}*{\mathcal{R}}_{j}f(t)-{\mathcal{R}}_{j}f(t)\|_{p}^{q}{\mathord{{\rm d}}}t=0.$
It must be noticed that for $q=\infty$ or $p=\infty$, the above property does not hold since $C_{b}({\mathbb{R}}^{d})$ is not dense in $L^{\infty}$. However, for $f\in{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{\alpha}_{\infty,\infty}$, by (2.4), it holds that for any $\alpha^{\prime}<\alpha$,
$\|f_{n}-f\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{\alpha^{\prime}}_{\infty,\infty}}\leqslant C\|f\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{\alpha^{\prime}}_{\infty,\infty}}n^{\alpha-\alpha^{\prime}}.$
This is also the only reason that we require $p,q<\infty$ below.
@@
Let $(X_{t})_{t\in[0,T]}$ be a ${\mathbb{R}}^{d}$-valued stochastic process on a probability space $(\Omega,{\mathscr{F}},{\mathbf{P}})$. For $\theta\in(0,1)$, one says that $(X,{\mathbb{B}})$ satisfies Krylov’s estimate with parameter $\theta$ if for any $f\in{\mathbb{L}}^{1}_{T}C_{b}\cap{\mathbb{B}}$, $m\in{\mathbb{N}}$, and any stopping times $\tau_{0}\leqslant\tau_{1}\leqslant T$ with $\tau_{1}-\tau_{0}\leqslant\delta$,
$\left\|\int^{\tau_{1}}_{\tau_{0}}f(s,X_{s}){\mathord{{\rm d}}}s\right\|_{L^{m}(\Omega)}\leqslant C\delta^{\frac{1+\theta}{2}}\|f\|_{\mathbb{B}},$
where the constant $C=C(m,{\mathbb{B}},d,T)>0$.
@@
Here we require $\theta>0$, that is crucial for Young’s integral in (4.5) below.
The following proposition is a direct conclusion of Krylov’s estimate.
@@
Suppose that $(X,{\mathbb{B}})$ satisfies Krylov’s estimate with parameter $\theta\in(0,1)$. For any $f\in{\mathbb{B}}$, let $f_{n}\in{\mathbb{L}}^{1}_{T}C_{b}\cap{\mathbb{B}}$ be such that
@@
Then the limit $\lim_{n\to\infty}\int^{\cdot}_{0}f_{n}(s,X_{s}){\mathord{{\rm d}}}s$ exists in $L^{p}(\Omega,C([0,T]))$ for any $p\geqslant 1$ and the limit is denoted by $A^{f}_{t}:=\lim_{n\to\infty}\int^{t}_{0}f_{n}(s,X_{s}){\mathord{{\rm d}}}s$, which does not depend on the choice of approximation sequence $(f_{n})_{n\in{\mathbb{N}}}$. Furthermore, for any $m\in{\mathbb{N}}$,
$\sup_{t\in[0,T]}\|A^{f}_{t}\|_{L^{m}(\Omega)}+\sup_{s\neq t\in[0,T]}\frac{\|A^{f}_{t}-A^{f}_{s}\|_{L^{m}(\Omega)}}{|t-s|^{(1+\theta)/2}}\lesssim_{C}\|f\|_{\mathbb{B}}.$ (4.2)
We now show a substitution formula for Young’s integrals that will be used to show the uniqueness by Zvonkin’s transformation.
@@
Suppose that $(X,{\mathbb{B}})$ satisfies Krylov’s estimate with parameter $\theta\in(0,1)$ and for any $m\in{\mathbb{N}}$,
$\|X_{0}\|_{L^{m}(\Omega)}+\|X_{t}-X_{s}\|_{L^{m}(\Omega)}\lesssim_{C}|t-s|^{1/2},\quad\forall s,t\in[0,T].$ (4.3)
@@
By (4.2), (4.3) and Kolmogorov’s continuity criterion, for any $\alpha\in(\frac{1}{2},\frac{1+\theta}{2})$, $\beta\in(0,\frac{1}{2})$ and $m\in{\mathbb{N}}$, there is a constant $C>0$ such that for all $f\in{\mathbb{B}}$,
$\big{\|}[A^{f}_{\cdot}]_{\alpha}\big{\|}_{L^{m}(\Omega)}\lesssim_{C}\|f\|_{\mathbb{B}},\ \ \big{\|}[X_{\cdot}]_{\beta}\big{\|}_{L^{m}(\Omega)}<\infty.$ (4.6)
Below we fix $\alpha>\frac{1}{2},\beta<\frac{1}{2}$ with $\alpha+\beta>1$ and a sample point $\omega$ such that
@@
Since $\alpha+\beta>1$, by the sewing lemma (cf. [17, Lemma 4.2]), we have
@@
Now, let $\{f_{n}\}_{n=1}^{\infty}\subset{\mathbb{L}}^{1}_{T}C_{b}\cap{\mathbb{B}}$ be such that $\lim_{n\to\infty}\|f_{n}-f\|_{{\mathbb{B}}}=0$. By definition, it is easy to see that
@@
$\theta_{0}:=1+\alpha_{b}-\tfrac{d}{p_{b}}-\tfrac{2}{q_{b}},\ \ b_{n}(t,x):=b(t,\cdot)*\phi_{n}(x).$
By the property of convolution, we have
$\begin{split}\kappa^{b_{n}}_{1}=\|b_{n}\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}}&\leqslant\|b\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}}=\kappa^{b}_{1},\\ \kappa^{b_{n}}_{2}=\|\mathord{{\rm div}}b_{n}\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{-2-\alpha_{b}}_{p_{b},q_{b}/(q_{b}-1)}}&\leqslant\|\mathord{{\rm div}}b\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{-2-\alpha_{b}}_{p_{b},q_{b}/(q_{b}-1)}}=\kappa^{b}_{2}.\end{split}$ (4.10)
We consider the following approximation SDE:
${\mathord{{\rm d}}}X^{n}_{t}=b_{n}(t,X^{n}_{t}){\mathord{{\rm d}}}t+\sqrt{2}{\mathord{{\rm d}}}W_{t},\ \ X^{n}_{0}\sim\mu\in{\mathcal{P}}({\mathbb{R}}^{d}).$ (4.11)
For simplicity of notations, we write
${\mathscr{I}}^{\alpha_{b}}_{p_{b},q_{b}}:=\Big{\{}(\alpha,p,q)\in[-1,0]\times[1,\infty)^{2}:\tfrac{q_{b}}{2}\leqslant q\leqslant q_{b},\ \ p\leqslant p_{b},\ \ \alpha-\tfrac{d}{p}-\tfrac{2}{q}\geqslant\alpha_{b}-\tfrac{d}{p_{b}}-\tfrac{2}{q_{b}}\Big{\}}.$
Now we can show the following crucial Krylov estimate.
Lemma 4.5.
For any $(\alpha,p,q)\in{\mathscr{I}}^{\alpha_{b}}_{p_{b},q_{b}}$, $(X^{n},{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q})$ satisfies Krylov’s estimate with parameter $\theta_{0}$ uniformly in $n$.
Proof.
Given $f\in C^{\infty}_{c}([0,T]\times{\mathbb{R}}^{d})$, let $u_{n}\in{\mathbb{L}}^{\infty}_{T}C^{\infty}_{b}$ be the unique smooth solution to the following backward PDE
$\partial_{t}u_{n}+\Delta u_{n}-\lambda u_{n}+b_{n}\cdot\nabla u_{n}+f=0,\ \ u_{n}(T)=0.$
For any $(\alpha,p,q)\in{\mathscr{I}}^{\alpha_{b}}_{p_{b},q_{b}}$, by (3.4) and (4.10), there is a $\lambda_{0}\geqslant 1$ so that for all $\theta\in(0,2+\alpha_{b}+\frac{d}{p_{b}}-\frac{2}{q_{b}}]$ and $\lambda\geqslant\lambda_{0}$,
By Itô’s formula, we have
$u_{n}(t,X^{n}_{t})=u_{n}(0,x)+\int^{t}_{0}(\lambda u_{n}-f)(s,X^{n}_{s}){\mathord{{\rm d}}}s+\sqrt{2}\int^{t}_{0}\nabla u_{n}(s,X^{n}_{s}){\mathord{{\rm d}}}W_{s}.$
$\left|\int^{\tau_{1}}_{\tau_{0}}f(s,X^{n}_{s}){\mathord{{\rm d}}}s\right|\leqslant 2\|u_{n}\|_{\infty}+\lambda\delta\|u_{n}\|_{\infty}+\sqrt{2}\left|\int^{\tau_{1}}_{\tau_{0}}\nabla u_{n}(s,X^{n}_{s}){\mathord{{\rm d}}}W_{s}\right|.$
By BDG’s inequality, we have for any $m\geqslant 2$,
$\left\|\int^{\tau_{1}}_{\tau_{0}}f(s,X^{n}_{s}){\mathord{{\rm d}}}s\right\|_{L^{m}(\Omega)}$ $\lesssim(1+\lambda\delta)\|u_{n}\|_{\infty}+m\left\|\left(\int^{\tau_{1}}_{\tau_{0}}|\nabla u_{n}(s,X^{n}_{s})|^{2}{\mathord{{\rm d}}}s\right)^{1/2}\right\|_{L^{m}(\Omega)}$
$\lesssim(1+\lambda\delta)\|u_{n}\|_{\infty}+m\delta^{1/2}\|\nabla u_{n}\|_{\infty}.$
by (4.12) and embedding (2.8), we have
and
Thus
$\left\|\int^{\tau_{1}}_{\tau_{0}}f(s,X^{n}_{s}){\mathord{{\rm d}}}s\right\|_{L^{m}(\Omega)}$ $\lesssim\left[(1+\lambda\delta)(1+\lambda)^{-\frac{1+\theta_{0}}{2}}+m\sqrt{\delta}(1+\lambda)^{-\frac{\theta_{0}}{2}}\right]\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q}}.$
In particular, if we take $\lambda=\lambda_{0}\vee\delta^{-1}$, then
$\left\|\int^{\tau_{1}}_{\tau_{0}}f(s,X^{n}_{s}){\mathord{{\rm d}}}s\right\|_{L^{m}(\Omega)}\lesssim m\delta^{\frac{1+\theta_{0}}{2}}\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q}}.$ (4.13)
By a standard approximation, the above estimate still holds for any $f\in{\mathbb{L}}^{1}_{T}C_{b}\cap{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q}$. The proof is complete. ∎
Remark 4.6.
By Stirling’s formula, estimate (4.13) also implies that for some $c_{0}>0$,
$\sup_{n}{\mathbf{E}}\exp\left\{c_{0}\left|\int^{T}_{0}f(s,X^{n}_{s}){\mathord{{\rm d}}}s\right|/\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q}}\right\}<\infty.$
@@
$\partial_{t}u^{\lambda}+\Delta u-\lambda u^{\lambda}+b\odot\nabla u^{\lambda}+\mathord{{\rm div}}b\preceq u^{\lambda}+b=0,\ \ u(T)=0,$
@@
$\|u^{\lambda}\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{2+\alpha_{b}-2/q_{b}-\theta}_{p_{b},1}}\lesssim(1+\lambda)^{-\frac{\theta}{2}}\|b\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}},$ (4.14)
@@
where $\Phi^{-1}$ stands for the inverse of $x\mapsto\Phi(t,x)$. Now we show the following crucial Zvonkin transformation.
@@
Let $(\mathfrak{F},X,W)$ be a weak solution of SDE (1.1) in the sense of Definition 1.1. Suppose that for any $(\alpha,p,q)\in{\mathscr{I}}^{\alpha_{b}}_{p_{b},q_{b}}$, $(X,{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q})$ satisfies Krylov’s estimate with parameter $\theta_{0}$ as above. Then $Y_{t}:=\Phi(t,X_{t})$ solves the following SDE:
@@
$\tilde{b}(t,y):=\lambda u(t,\Phi^{-1}(t,y)),\ \ \tilde{\sigma}(t,y):=\sqrt{2}\nabla\Phi^{-1}(t,\Phi^{-1}(t,y)).$
@@
$|\tilde{b}(t,y)-\tilde{b}(t,y^{\prime})|\leqslant C|y-y^{\prime}|,\ \ |\tilde{\sigma}(t,y)-\tilde{\sigma}(t,y^{\prime})|\leqslant C|y-y^{\prime}|^{\gamma},$ (4.18)
and for some $c_{0}>0$,
$|\tilde{\sigma}(t,y)\xi|^{2}\geqslant c_{0}|\xi|^{2},\ \ \xi\in{\mathbb{R}}^{d}.$ (4.19)

### v3

@@ 3.2. Supercritical case: $\frac{d}{p}+\frac{2}{q}<2+\alpha$
By [51, Theorem 2.2], there is a constant $C=C(\Theta,\alpha,p,q)>0$ such that for all $n\in{\mathbb{N}}$,
$\|u_{n}\|_{{\mathbb{L}}^{\infty}_{T}}+\|u_{n}\|_{{\mathscr{V}}_{T}}\leqslant C\|f_{n}\|_{{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}}\leqslant C\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}}.$ (3.14)
By (2.9) and (2.10), it is easy to see that for $\frac{1}{p_{0}}:=\frac{1}{p}_{b}+\frac{1}{2}\leqslant\frac{1}{d}+\frac{1}{2}$,
$\|fg\|_{{\mathbf{H}}^{-2}_{2}}\lesssim\|fg\|_{{\mathbf{H}}^{-1}_{p_{0}}}\leqslant\|f\|_{{\mathbf{H}}^{-1}_{p_{b}}}\|g\|_{{\mathbf{H}}^{1}_{2}}.$
Thus by (3.14), we have for Lebesgue almost all $t$,
$\|\partial_{t}u_{n}(t)\|_{{\mathbf{H}}^{-3}_{2}}$ $\leqslant\|\Delta u_{n}(t)+\mathord{{\rm div}}((b_{n}u_{n})(t))-\mathord{{\rm div}}b_{n}(t)\,u_{n}(t)+f_{n}(t)\|_{{\mathbf{H}}^{-3}_{2}}$
$\lesssim\|u_{n}(t)\|_{2}+\|(b_{n}u_{n})(t)\|_{{\mathbf{H}}^{-2}_{2}}+\|\mathord{{\rm div}}b_{n}(t)u_{n}(t)\|_{2}+\|f_{n}(t)\|_{{\mathbf{H}}^{\alpha}_{p}}$
$\lesssim\|u_{n}\|_{{\mathbb{L}}^{\infty}_{T}L^{2}}+\|b_{n}(t)\|_{{\mathbf{H}}^{-1}_{p_{b}}}\|u_{n}(t)\|_{{\mathbf{H}}^{1}_{2}}+\|\mathord{{\rm div}}b_{n}(t)\|_{\infty}\|u_{n}(t)\|_{2}+\|f(t)\|_{{\mathbf{H}}^{\alpha}_{p}}$
$\lesssim\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}}+\|b(t)\|_{{\mathbf{H}}^{-1}_{p_{b}}}\|u_{n}(t)\|_{{\mathbf{H}}^{1}_{2}}+\|\mathord{{\rm div}}b(t)\|_{\infty}\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}}+\|f(t)\|_{{\mathbf{H}}^{\alpha}_{p}},$
where the implicit constant does not depend on $n$ and $t$. Hence, for all $0\leqslant t_{0}<t_{1}\leqslant T$,
$\sup_{n}\|u_{n}(t_{1})-u_{n}(t_{0})\|_{{\mathbf{H}}^{-3}_{2}}\leqslant\sup_{n}\int^{t_{1}}_{t_{0}}\|\partial_{t}u_{n}(t)\|_{{\mathbf{H}}^{-3}_{2}}{\mathord{{\rm d}}}t$
$\quad\lesssim(t_{1}-t_{0})+\left(\int^{t_{1}}_{t_{0}}\|b(t)\|^{2}_{{\mathbf{H}}^{-1}_{p_{b}}}{\mathord{{\rm d}}}t\right)^{1/2}+\int^{t_{1}}_{t_{0}}\Big{(}\|\mathord{{\rm div}}b(t)\|_{\infty}+\|f(t)\|_{{\mathbf{H}}^{\alpha}_{p}}\Big{)}{\mathord{{\rm d}}}t,$
which in turn implies that
$\lim_{|t_{1}-t_{0}|\to 0}\sup_{n}\|u_{n}(t_{1})-u_{n}(t_{0})\|_{{\mathbf{H}}^{-3}_{2}}=0.$
By Aubin-Lions’ lemma (see [39, Corollary 4] for instance) and the above uniform continuity, there is a function $u\in{\mathbb{L}}^{\infty}_{T}\cap{\mathscr{V}}_{T}$ and subsequence $\{n_{k}\}$ such that for each $t$ and $\varphi\in C^{\infty}_{c}({\mathbb{R}}^{d})$,
@@
$\|\langle(b_{n_{k}}-b)u_{n_{k}},\nabla\varphi\rangle\|_{L^{1}_{T}}\leqslant C\|(b_{n_{k}}-b)\chi\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{H}}^{\alpha_{b}}_{p_{b}}}\|u_{n_{k}}\|_{{\mathscr{V}}_{T}}\|\nabla\varphi\|_{{\mathscr{V}}_{T}}\stackrel{{\scriptstyle k\to\infty}}{{\to}}0,$
@@
Hence, by $\mathord{{\rm div}}b\in L^{\infty}$ and commutating the order of limits (guaranteed by the above two limits),
$\lim_{n\to\infty}\langle u,b\cdot\nabla u_{n}\rangle=\lim_{n\to\infty}\lim_{m\to\infty}\langle u_{m},b\cdot\nabla u_{n}\rangle$ $=-\lim_{n\to\infty}\lim_{m\to\infty}\langle\mathord{{\rm div}}b\,u_{m},u_{n}\rangle-\lim_{n\to\infty}\lim_{m\to\infty}\langle b\cdot\nabla u_{m},u_{n}\rangle$
$=-\langle\mathord{{\rm div}}b\,u,u\rangle-\lim_{m\to\infty}\langle b\cdot\nabla u_{m},u\rangle,$
@@
$\lim_{n\to\infty}\langle u,b\cdot\nabla u_{n}\rangle=-\langle\mathord{{\rm div}}b\,u,u\rangle/2.$
@@
In this section, we establish the weak well-posedness of SDE (1.1) under the subcritical assumption. To achieve this, we employ the classical Zvonkin transformation (see [50, Section 5]), ensuring both the existence and uniqueness of a weak solution.
@@
  • For any bounded measurable function $g:[0,T]\times{\mathbb{R}}^{d}\to{\mathbb{R}}$ such that the function $x\to g(t,x)$ is uniformly Lipschitz with respect to $t\in[0,T]$, the product $gf$ is well-defined for all $f\in{\mathbb{B}}$, and there is a constant $C_{g}>0$ such that for all $f\in{\mathbb{B}}$,
@@
$\lim_{n\to\infty}\|f_{n}-f\|^{q}_{{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q}}\leqslant\int^{T}_{0}\sum_{j\geqslant-1}2^{\alpha jq}\lim_{n\to\infty}\|\phi_{n}*{\mathcal{R}}_{j}f(t)-{\mathcal{R}}_{j}f(t)\|_{p}^{q}{\mathord{{\rm d}}}t=0.$ (4.2)
Throughout most of this section, we limit our analysis to the case $p,q<\infty$ when considering the space ${\mathbb{B}}={\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q}$. The case $p=\infty$ or $q=\infty$ will be addressed later in the proof of Theorem 1.2.
@@
Let $(X_{t})_{t\in[0,T]}$ be a ${\mathbb{R}}^{d}$-valued stochastic process on a probability space $(\Omega,{\mathscr{F}},{\mathbf{P}})$. For $\theta\in(0,1)$ and $m>2$, one says that $(X,{\mathbb{B}})$ satisfies Krylov’s estimate with parameters $(m,\theta)$ and constant $C>0$ if for any $f\in{\mathbb{L}}^{1}_{T}C_{b}\cap{\mathbb{B}}$ and any stopping times $0\leqslant\tau_{0}\leqslant\tau_{1}\leqslant T$ with $\tau_{1}-\tau_{0}\leqslant\delta$,
$\left\|\int^{\tau_{1}}_{\tau_{0}}f(s,X_{s}){\mathord{{\rm d}}}s\right\|_{L^{m}(\Omega)}\leqslant C\delta^{\frac{1+\theta}{2}}\|f\|_{\mathbb{B}}.$ (4.3)
@@
Here we require $\theta>0$, which is crucial for Young’s integral in (4.7) below.
The following proposition is a direct conclusion of Krylov’s estimate and the Kolmogorov continuity theorem.
@@
Suppose that $(X,{\mathbb{B}})$ satisfies Krylov’s estimate with parameter $(m,\theta)\in(2,\infty)\times(0,1)$. For any $f\in{\mathbb{B}}$, let $f_{n}\in{\mathbb{L}}^{1}_{T}C_{b}\cap{\mathbb{B}}$ be such that
@@
Then the limit $A^{f}_{\cdot}:=\lim_{n\to\infty}\int^{\cdot}_{0}f_{n}(s,X_{s}){\mathord{{\rm d}}}s$ exists in $L^{m}(\Omega,C([0,T]))$ and does not depend on the choice of approximation sequence. Furthermore,
$\sup_{t\in[0,T]}\|A^{f}_{t}\|_{L^{m}(\Omega)}+\sup_{s\neq t\in[0,T]}\frac{\|A^{f}_{t}-A^{f}_{s}\|_{L^{m}(\Omega)}}{|t-s|^{(1+\theta)/2}}\leqslant C\|f\|_{\mathbb{B}}.$ (4.4)
Next we show a substitution formula for Young’s integrals that will be used to show the uniqueness by Zvonkin’s transformation.
@@
Let $m>2$ and $\theta\in(0,1)$ satisfy $\theta m>2$. Suppose that $(X,{\mathbb{B}})$ satisfies Krylov’s estimate with parameters $(m,\theta)$, and for any $\beta\in(0,\frac{1}{2})$,
$\left\|\sup_{s\not=t\in[0,T]}|X_{t}-X_{s}|/|t-s|^{\beta}\right\|_{L^{m}(\Omega)}<\infty.$ (4.5)
@@
Since $\theta m>2$, one can choose $\alpha\in(\frac{1}{2},\frac{1+\theta}{2}-\frac{1}{m})$ and $\beta\in(0,\frac{1}{2})$ so that
$\alpha+\beta>1.$
By (4.4), (4.5) and Kolmogorov’s continuity criterion, there is a constant $C>0$ such that for all $f\in{\mathbb{B}}$,
$\big{\|}[A^{f}_{\cdot}]_{\alpha}\big{\|}_{L^{m}(\Omega)}\leqslant C\|f\|_{\mathbb{B}},\ \ \big{\|}[X_{\cdot}]_{\beta}\big{\|}_{L^{m}(\Omega)}<\infty.$ (4.8)
Below we fix a sample point $\omega$ such that
@@
Since $\alpha+\beta>1$, by the sewing lemma (cf. [16, Lemma 4.2]), we have
@@
Now, let $\{f_{n}\}_{n=1}^{\infty}\subset{\mathbb{L}}^{1}_{T}C_{b}\cap{\mathbb{B}}$ be such that $\lim_{n\to\infty}\|f_{n}-f\|_{{\mathbb{B}}}=0$. By definition, one sees that
@@
Below, we always suppose that $b=b_{1}+b_{2}$ satisfies
  • ($\widetilde{\bf H}^{\rm sub}_{b}$)
@@
$\partial_{t}u^{\lambda}+\Delta u-\lambda u^{\lambda}+{b_{1}}\odot\nabla u^{\lambda}+\mathord{{\rm div}}{b_{1}}\preceq u^{\lambda}+{b_{1}}=0,\ \ u(T)=0,$
@@
$\|u^{\lambda}\|_{{\mathbb{L}}^{\infty}_{T}{\mathbf{B}}^{2+\alpha_{b}-2/q_{b}-\theta}_{p_{b},1}}\lesssim(1+\lambda)^{-\frac{\theta}{2}}\|{b_{1}}\|_{{\mathbb{L}}^{q_{b}}_{T}{\mathbf{B}}^{\alpha_{b}}_{p_{b},q_{b}}},$ (4.12)
@@
where $\Phi^{-1}$ stands for the inverse of $x\mapsto\Phi(t,x)$.
Below, for simplicity of notations, we write
${\mathbf{I}}^{\alpha_{b}}_{p_{b},q_{b}}:=\Big{\{}(\alpha,p,q)\in[-1,0]\times[1,\infty)^{2}:\tfrac{q_{b}}{2}\leqslant q\leqslant q_{b},\ \ p\leqslant p_{b},\ \ \alpha-\tfrac{d}{p}-\tfrac{2}{q}\geqslant\alpha_{b}-\tfrac{d}{p_{b}}-\tfrac{2}{q_{b}}\Big{\}}.$
We now proceed to show the following crucial Zvonkin transformation.
@@
Let $m>2$ and $\theta\in(0,1)$ satisfy $\theta m>2$. Let $(\mathfrak{F},X,W)$ be a weak solution of SDE (1.1) in the sense of Definition 1.1. Suppose that for any $(\alpha,p,q)\in{\mathbf{I}}^{\alpha_{b}}_{p_{b},q_{b}}$, $(X,{\mathbb{L}}^{q}_{T}{\mathbf{B}}^{\alpha}_{p,q})$ satisfies Krylov’s estimate with parameter $(m,\theta)$. Then $Y_{t}:=\Phi(t,X_{t})$ solves the following SDE:
@@
$\tilde{b}(t,y):=\lambda u(t,\Phi^{-1}(t,y))+(b_{2}\cdot\nabla\Phi)(t,\Phi^{-1}(t,y)),\ \ \tilde{\sigma}(t,y):=\sqrt{2}\nabla\Phi(t,\Phi^{-1}(t,y)).$
@@
${|\tilde{b}(t,y)|\leqslant C(\lambda+c_{0}+c_{1}|y|)},\ \ |\tilde{\sigma}(t,y)-\tilde{\sigma}(t,y^{\prime})|\leqslant C|y-y^{\prime}|^{\gamma},$ (4.17)
where $C=C(T,\gamma,\Theta)$, and for all $\xi\in{\mathbb{R}}^{d}$,
$|\xi|/8\leqslant|\tilde{\sigma}(t,y)\xi|\leqslant 8|\xi|.$ (4.18)